OpenStudy (anonymous):

5-h=2gk+4 solve for g

OpenStudy (luigi0210):

Hello @TLCarlisle41 Welcome to openstudy :)

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

No

OpenStudy (luigi0210):

Well start off by moving that 4 over to the other side

OpenStudy (anonymous):

So I subtract 4 from both sides

OpenStudy (luigi0210):

Yup, and what do you get?

OpenStudy (anonymous):

1h or h

OpenStudy (anonymous):

1-h

OpenStudy (luigi0210):

You don't combine numbers and terms if they're not alike

OpenStudy (anonymous):

Is it 1-h

OpenStudy (luigi0210):

Yup, so now you have: 1-h=2gk

OpenStudy (luigi0210):

The 2 and k are being multiplied so how do you undo multiplication?

OpenStudy (anonymous):

Divide

OpenStudy (luigi0210):

Yea, so just divide everything by 2k

OpenStudy (anonymous):

Does the h and k cancel each other out?that would leave1=2g

OpenStudy (luigi0210):

No, you can't cancel

OpenStudy (anonymous):

So it would be 1-h/2k=2g/2k

OpenStudy (anonymous):

Thanks

OpenStudy (luigi0210):

I'm sorry about that, but no it would be: \[\frac{1-h}{2k}=g\]